# LCLs beyond bounded degrees - Locally Finite Labelings

### v1

@@ 1.3 A Tour of the Gap Results
Seeing as the proof of the gap results is quite long and technical, we provide an overview of how the proofs work.
@@
For LCLs in bounded degree trees, there is a beautiful half page argument, that any LCL can be expressed in the node edge checkable formalism and vice versa. The key challenge here is that, in an LCL, correctness depends on the entire $r$ hop neighborhood of a node, while in the case of the node-edge checkable formalism the correctness must be checked by looking only at adjacent halfedges. Phrased in an imprecise way, the checking radius in the node-edge checkable formalism is $0.5$.
@@
Initially, it seems like extending this approach to LFLs is straightforward, there is still just a finite set of allowed configurations $\mathcal{C}$. However, crucially, the $r$-hop neighborhood of a node $v$ is no longer isomorphic to one of these configurations $C\in\mathcal{C}$.
@@
Because LFLs in the node–edge checkable formalism have an effective checking radius of $0.5$, the labeling feasibility on the opposite side of the cut depends only on the label chosen for $e_{1}$, and not on the internal structure of $T_{1}$. This locality allows us to classify trees solely by the subsets of labels they permit on their boundary halfedges.
@@
We then extend this idea from single edges to paths. Instead of cutting the tree at one edge, we cut out a path $P$ by separating it from the rest of the tree at two edges. Let $H$ denote the induced subgraph consisting of $P$ and all components attached to it, and let $e_{s}$ and $e_{t}$ be the two remaining boundary halfedges of $H$. We now ask: which pairs of labels can be assigned to $e_{s}$ and $e_{t}$ such that the labeling of $H$ can be completed? Going one step further, we ask whether there exist sets $X_{s},X_{t}\subseteq\Sigma_{\text{out}}$ such that every choice of labels from $X_{s}$ and $X_{t}$ can be extended to a valid labeling of $H$.
Figure 5: By cutting out a piece of the graph at the ends of some path $P$, splits the tree into three parts $T_{s},H,T_{t}$. The parts are connected by the edges $e_{s}$ and $e_{t}$.
@@
We say that $X_{s},X_{t}$ form an independent class for $H$, since the choices of labels from $X_{s}$ and $X_{t}$ can be made independently. We study the existence of such independent classes using the notion of types of trees introduced earlier. As before, the finiteness of the output alphabet $\Sigma_{\text{out}}$ implies that there are only finitely many possible types into which subgraphs $H$ can fall.
@@
We can use these longer versions together with a fast algorithm $\mathcal{A}$ in the following way. To simulate $\mathcal{A}$ on $H$, we first replace it by a pumped version $H^{\prime}$ with identical boundary behavior, chosen long enough so that when simulating $\mathcal{A}$ at the midpoint of $H^{\prime}$, the algorithm cannot see outside of $H^{\prime}$. Since $\mathcal{A}$ must be correct, the labeling it produces induces a non-empty independent class for $H^{\prime}$, and by extension for $H$. In this way, fast algorithms give rise to independent classes.
@@
The final step is to show that if a problem admits a deterministic algorithm with complexity $o(n^{1/k})$, then this recursive decoupling must indeed be possible $k$ times. When pumping up a path for the first time, it results in a path of length $\Omega(n^{1/k})$, but when doing it for the second time, each node of that path might already have a pumped path of length $\Omega(n^{1/k})$ attached to it, so the resulting instance needs $\Omega(n^{2/k})$ many nodes. If we tried to do this construction for more than $k$ steps, we would need more than $n$ nodes and so the simulation argument does not work anymore. Hence, an algorithm with complexity $o(n^{1/k})$ can be used to repeat the above process $k$ times. By then abusing the fact that this decoupling can be done $k$ times, we obtain an algorithm with complexity $O(n^{1/(k+1)})$.
@@ 2 Preliminaries
3 LCLs with infinitely many Configurations
In this section we show that if our problem definitions are allowed to make infinitely many case distinctions, we can design problems of arbitrary polynomial complexities in unbounded degree trees. We will build on the intuition given in the introduction and explain how we generalise this idea to achieve complexities $\Theta(n^{r})$, for any rational $0<r\leq\frac{1}{2}$. Note that Theorem 1 works for any $0<r\leq 1$, the more general proof, which can also acheive $\frac{1}{2}<r\leq 1$ is deferred to Appendix B.
We want to retain the Local of Locally Checkable Loabelings in the sense that solutions depend only on the local information around a node. To this end, we call a problem $r$-checkable if
  1. 1.
  2. 2.
With this we can state Theorem 1 again. See 1 Keep in mind, that allowing for an unbounded number of lables (as would be necessary for e.g. $\Delta+1$ coloring) immediately will also lead to an unbounded number of configurations. So in some sense the result of this section also shows the necessity to restrict the number of labels for our definition of LFLs.
To prove Theorem 1, we design a problem that is hard only if it contains a path with a strictly increasing number of leaves attached to each node of the path. By scaling how fast this number of nodes grows, we can change the complexity of our problem. In full generality, we would just take any strictly positive injective function $f$, then we require the starting node to have exactly $f(1)$ leaves attached, the second node in the path to have $f(2)$, the third $f(3)$ and so on.
Instead we will stick to just nicely behaved polynomials of the form $f(x)=x^{\alpha}$, for some constant $\alpha$.
Definition 4 ($\alpha$-path).
Let $\alpha\geq 1$. In any tree $T=(V,E)$, a path $P=(v_{1},v_{2},\ldots,v_{L})$, of length $L\geq 2$, is an $\alpha$-path, if
  1. 1.
    $v_{1}$ has only a single leaf-neighbor and $v_{2}$ has exactly $\lfloor 2^{\alpha}\rfloor$ leaf-neighbors.
  2. 2.
    $v_{i}$ has exactly $\lfloor i^{\alpha}\rfloor$ leaf-neighbors and a neighbor $v_{i-1}$ with exactly $\lfloor(i-1)^{\alpha}\rfloor$ leaf neighbors, for all $1<i\leq L$.
To simplify notation a little bit, we fix $f(x)=\lfloor x^{\alpha}\rfloor$, so the starting condition 1 now says that the node itself has to have $f(1)=1$ many leaf neighbors and $v_{i}$ must have $f(i)$ leaf neighbors.
Notice that for $\alpha=1$, we exactly replicate the problem described in the introduction ($v_{i}$ has to have $i$ leaves attached).
The Family of Problems:
We now formulate the problem as a $3$-checkable problem. Each node must output 1 if it belongs to an $\alpha$-path, and 0 otherwise. To enforce this, we encode the defining conditions of an $\alpha$-path into local configurations. These conditions depend only on (i) the number of leaves attached to a node $v$, and (ii) the number of leaves attached to $v$’s non-leaf neighbors.
Determining the number of leaves adjacent to $v$ requires only distance-2 information. However, verifying the number of leaves adjacent to $v$’s neighbors requires access to distance-3 information [^4] Remember, that wether or not two nodes at distance 2 have an edge connecting them is distance 3 information based on our definitions.. Thus, the checking radius of our LCL is $r=3$. Importantly, the $\alpha$-path constraints do not uniquely specify the entire 3-hop neighborhood, so the remaining structure within that radius can be arbitrary.
The problem uses only labels 0,1{0,1}: nodes on an $\alpha$-path must output 1, and all others output 0. To define the problem, it suffices to describe the 2-hop neighborhoods in which nodes are required to output 1. All other neighborhoods force output 0. There are exactly two types of local configurations that trigger output 1, shown in Figure 6: one that initiates a path (right) and one that propagates the label 1 along an $\alpha$-path (left). Note that each configuration represents an infinite family of possible 2-hop neighborhoods.
Figure 6: Key local configurations of $\Pi(\alpha)$, with center nodes indicated by arrows. Nodes without an explicitly specified output may take an arbitrary label (0 or 1). Leaves are represented only by downward edges, while the doted nodes stand for arbitrarily many (potentially 0) non-leaf neighbors. Note the outermost edges in both configurations may also connect to arbitrary, possibly non-leaf, nodes.
     Right : Start of an $\alpha$-path. The center node must output 1 if it has exactly one leaf neighbor, while at least one other non-leaf neighbor has exactly $f(2)$ leaves attached.
     Left: Propagation within an $\alpha$-path. The center node has $f(i)$ leaves (for some $i\in\mathbb{N}$), and one of its neighbors must have f(i1)f(i-1) leaves attached and output 1. So the configuration essentially requires a predecessor. This configuration also acts as the end of an $\alpha$-path, since the dotted neighbors on the left must not necessarily continue the chain.
Observation 5.
In an instance of $\Pi(\alpha)$ a node has to output 1 if and only if it is a node of an $\alpha$-path.
Proof.
Let $P=(v_{1},v_{2},\ldots,v_{L})$ be an $\alpha$-path. By definition, $v_{1}$ has exactly one leaf neighbor and a second neighbor $v_{2}$ with $f(2)$ leaf neighbors. Hence $v_{1}$ matches the right configuration of Figure 6 and must output $1$. Now suppose inductively that $v_{i-1}$ outputs $1$ for some $1<i\leq L$. Then $v_{i}$ has $f(i)$ leaf neighbors and a neighbor $v_{i-1}$ with f(i1)f(i-1) leaves and output $1$. Thus $v_{i}$ matches one of the left configurations of Figure 6 and must also output $1$. By induction, every node $v_{i}$ of $P$ outputs $1$.
For the converse, assume a node $v$ outputs $1$. Then $v$ must match one of the two types of configurations.
  • If $v$ matches the right configuration, then by definition it has exactly one leaf neighbor and one non-leaf neighbor $v_{2}$ with $f(2)$ leaf neighbors. This already certifies the existence of an $\alpha$-path $(v,v_{2})$ of length $2$.
  • If $v$ matches a left configuration corresponding to some index $i$, then $v$ has $f(i)$ leaf neighbors, and it must have a neighbor $v_{i-1}$ with f(i1)f(i-1) leaves that also outputs $1$, this neighbor must in turn have a neighbor $v_{i-2}$ with f(i2)f(i-2) neighbors. Repeating this argument recursively, we obtain a strictly decreasing sequence of leaf counts until eventually reaching a node that matches the starting configuration. Tracing back this chain yields a valid $\alpha$-path that contains $v$.
Hence a node outputs $1$ if and only if it is part of an $\alpha$-path. ∎
The only important property that is relevant for the distributed complexity of $\Pi(\alpha)$ is, how long these $\alpha$-paths are
Lemma 6.
In any instance of $\Pi(\alpha)$ on an $n$ node graph. The maximum length of an $\alpha$-path is at most $O(n^{1/(\alpha+1)})$. Furthermore, there exists an $\alpha$-path of length $\Omega(n^{1/(\alpha+1)})$ using less than $n$ nodes.
Proof.
The definition of the problem restricts the paths of nodes that output 1 to be exactly those paths with the correct (increasing) sequence of leaves attached. It starts with a node that has $f(1)=1^{\alpha}=1$ leaves attached, then one with $f(2)$ leaves attached and so on. For a path to have length $L$, the number of attached leaves is
$\sum_{x=1}^{L}f(x)=\sum_{x=1}^{L}\lfloor x^{\alpha}\rfloor\leq\sum_{x=1}^{L}x^{\alpha}\leq\int_{0}^{L}x^{\alpha}\,dx=\frac{1}{\alpha}L^{\alpha+1}$
Hence the length $L$ of such a path is limited by the fact, that this number must be smaller than the total number of nodes $n$
$\frac{1}{\alpha}L^{\alpha+1}\leq n\iff L\leq(\alpha n)^{1/(\alpha+1)}\iff L\in O(n^{1/(\alpha+1)})$
Note that the total number of nodes required for a path of length $L$ is just the $L$ nodes of the path plus these attached leaves. So there exists an $L\in\Omega(n^{1/(\alpha+1)})$, such that an $\alpha$-path of length $L$ has less than $n$ nodes. ∎
This is enough to prove half of our main theorem.
Lemma 7.
For any real $0<b\leq\tfrac{1}{2}$, there exists a $3$-checkable problem using only labels $\{0,1\}$ with distributed complexity $\Theta(n^{b})$ in unbounded-degree trees.
Proof.
Since $b\leq\tfrac{1}{2}$, we can write $b=\tfrac{1}{1+\alpha}$ for some $\alpha\geq 1$. We argue that the problem $\Pi(\alpha)$ has distributed complexity
$\Theta\!\bigl(n^{1/(\alpha+1)}\bigr)=\Theta(n^{b}).$
We provide an $O(n^{1/(\alpha+1)})$-round deterministic algorithm and then show that any randomized algorithm requires at least $\Omega(n^{1/(\alpha+1)})$ rounds to succeed with probability greater than $1/2$. Both bounds follow from standard distributed arguments and Lemma 6.
Upper bound. Consider the trivial deterministic algorithm in which each node explores its $O(n^{1/(\alpha+1)})$-hop neighborhood and checks whether it is part of an $\alpha$-path. If so, it outputs $1$; otherwise, it outputs 0. By Lemma 6, the maximum length of any $\alpha$-path is $O(n^{1/(\alpha+1)})$, so each node can determine its membership correctly within that radius. Hence the problem is solvable in $O(n^{1/(\alpha+1)})$ rounds in the LOCAL model.
Lower bound. For the lower bound, we use a standard indistinguishability argument. Suppose a randomized LOCAL algorithm runs in time $T(n)$. Then for any two instances in which a node $v$ has the same $T(n)$-neighborhood, $v$’s output distribution must be identical in both.
Assume for contradiction that there exists a constant $\varepsilon>0$ such that $T(n)\leq\varepsilon n^{1/(\alpha+1)}$ and the algorithm succeeds with probability greater than $1/2$. By Lemma 6, for sufficiently large $n$ there exists an $\alpha$-path
$P=(v_{1},v_{2},\ldots,v_{L})$
of length $L>\varepsilon n^{1/(\alpha+1)}$ using fewer than $n$ nodes. We can pad $P$ to an $n$-node tree by, for example, attaching an additional path to $v_{L}$, which preserves the $\alpha$-path property.
Now define two instances:
  • $G_{1}$: the graph containing the full $\alpha$-path $P$.
  • $G_{2}$: the same graph with $v_{1}$ removed. Then $G_{2}$ contains no valid $\alpha$-path.
In $G_{1}$, the node $v_{L}$ must output $1$, while in $G_{2}$ the same node must output 0. However, since $L>\varepsilon n^{1/(\alpha+1)}\geq T(n)$, the $T(n)$-neighborhood of $v_{L}$ is identical in both graphs. Thus $v_{L}$ follows the same output distribution in $G_{1}$ and $G_{2}$. Consequently, with probability at least $1/2$, the algorithm errs on one of the two instances.
Hence the randomized complexity of $\Pi(\alpha)$ is $\Omega(n^{1/(\alpha+1)})$, which matches the deterministic upper bound.
Combining both bounds, the distributed complexity of $\Pi(\alpha)$ is $\Theta(n^{1/(\alpha+1)})=\Theta(n^{b})$, as claimed. ∎
To obtain complexities between $\omega(\sqrt{n})$ and $o(n)$, a more intricate construction is required. This involves less natural partition-gadgets and additional technical steps. Since we consider the statement of Lemma 7 already sufficient to motivate our definition of Locally Finite Labelings, we relegate this more elaborate construction to Appendix B, for readers interested in the full generality of the result.
@@ 4 Locally Finite Labelings
Examples of MIS and 3-coloring as LFLs were already given in the introduction, but we additionally provide the problem of 3-cycle detection, as a more involved example in Appendix A.
@@
Additionally we provide one nontrivial extension to LFLs that we will use to make our proofs much more elegant. We call this extension partially ordered LFLs ($\mathbf{poLFL}$). In partially ordered LFLs we allow wildcard labels in the configurations. For example, we could then describe the problem of coloring using configurations where the center node has color e.g. Red and it has a single optional neighbor with label Not Red. Now any neighbor that does not output color Red can be mapped onto this optional neighbor. We illustrate how to encode 3 coloring in just 3 intuitive configurations this way in Figure 7. This formalism is not just nice in the sense, that it gives us much more freedom to express problems in a concise and intuitive way, but it also makes some of our more involved proofs more manageable.
Figure 7: Using the ”not __” labels ($\{\bar{R},\bar{G},\bar{B}\}$) we can encode 3-coloring in the very intuitive way, where a node that outputs $R$ must have only neighbors that are ”not RR\,”.
@@
4.1 LFLs in trees - the Node Edge Checkable formalism
@@ A Locally Finite Labeling in the node–edge checkable formalism is a tuple
  • $\Sigma_{\text{out}}$ is a set of output labels, and we define $\Sigma_{\text{out}}^{*}=\{a,a^{*}\mid a\in\Sigma_{\text{out}}\}$. Thus for every output label $a$ we include a designated “relaxed” version $a^{*}$.
@@
Intuitively, each node $v$ inspects the multiset of labeled halfedges around it and checks whether some node constraint matches. If a pair $(a,b)$ appears exactly as often as required in a configuration $C$, the constraint is satisfied; if it appears more often, this is permitted only when $C$ contains $(a,b^{*})$, which serves as a wildcard allowing arbitrarily many occurrences.
MIS as a Node Edge Checkable Problem
Using no input labels and outputlabels $\left\{M,O,P\right\}$. The problem is given in the following way:
@@ 4.2 The TwigLFL
Based on this observation, we add exactly $m_{\ell}(X)$ required edges leading to nodes with labeling $\ell$ to $C_{X}$. Additionally, we add an optional edge leading to a node labeled $\ell$ if and only if every configuration $C\in X$ with $m_{\ell}(C)=m_{\ell}(X)$ contains such an optional edge. By 13, this condition implies that in fact all configurations in $X$ contain such an optional edge.
@@
If $N$ contains more than $m_{\ell}(X)$ such nodes, then by 13 every configuration in $X$ must contain an optional edge labeled $\ell$, and hence $C_{X}$ contains exactly one such optional edge. We map all remaining $\ell$-labeled nodes of $N$ to this optional edge.

### v2

@@ 1.3 A Tour of the Gap Results
The proof of the gap results is quite long and technical, we provide an overview of the main ideas.
@@
For LCLs in bounded degree trees, there is a beautiful half page argument, that any LCL can be expressed in the node-edge checkable formalism and vice versa. The key challenge here is that, in an LCL, correctness depends on the entire $r$ hop neighborhood of a node, while in the case of the node-edge checkable formalism the correctness must be checked by looking only at adjacent halfedges. Phrased in an imprecise way, the checking radius in the node-edge checkable formalism is $0.5$.
@@
Initially, it seems that extending this approach to LFLs is straightforward, there is still just a finite set of allowed configurations $\mathcal{C}$. However, crucially, the $r$-hop neighborhood of a node $v$ is no longer isomorphic to one of these configurations $C\in\mathcal{C}$.
@@
Because LFLs in the node–edge checkable formalism have an effective checking radius of $0.5$, we can classify trees solely by the subsets of labels they permit on their boundary halfedges.
@@
We then extend this idea from single edges to paths. Instead of cutting the tree at one edge, we cut out a path $P$ by separating it from the rest of the tree at two edges. Let $H$ denote the induced subgraph consisting of $P$ and all components attached to it, and let $e_{s}$ and $e_{t}$ be the two remaining boundary halfedges of $H$. This is depicted in Figure 5. We now ask: which pairs of labels can be assigned to $e_{s}$ and $e_{t}$ such that the labeling of $H$ can be completed? Going one step further, we ask whether there exist sets $X_{s},X_{t}\subseteq\Sigma_{\text{out}}$ such that we can draw labels for $e_{s}$ from $X_{s}$ and $e_{t}$ from $X_{t}$ independently and can then still obtain a valid labeling of $H$ for any two choices.
Figure 5: By cutting out a piece of the graph at the ends of some path $P$, splits the tree into three parts $T_{s},H,T_{t}$. The parts are connected by the edges $e_{s}$ and $e_{t}$. If we could pick labels for $e_{s}$ and $e_{t}$ without coordination, we could find labelings for $T_{s}$ and $T_{t}$ without the need for communication.
@@
We say that $(X_{s},X_{t})$ is an independent class for $H$, since the choices of labels from $X_{s}$ and $X_{t}$ can be made independently. We study the existence of such independent classes using the notion of types of trees introduced earlier. As before, the finiteness of the output alphabet $\Sigma_{\text{out}}$ implies that there are only finitely many possible types into which such subgraphs $H$ can fall.
@@
We can use these longer versions together with a fast algorithm $\mathcal{A}$ in the following way. To simulate $\mathcal{A}$ on $H$, we first replace it by a pumped version $H^{\prime}$ with identical boundary behavior, chosen long enough so that when simulating $\mathcal{A}$ at the midpoint of $H^{\prime}$, the algorithm cannot see outside of $H^{\prime}$. Since $\mathcal{A}$ must be correct, the labeling it produces in the middle of $H^{\prime}$ induces a non-empty independent class for $H^{\prime}$, and by extension for $H$. In this way, fast algorithms give rise to independent classes.
@@
The final step is to show that if a problem admits a deterministic algorithm with complexity $o(n^{1/k})$, then this recursive decoupling must indeed be possible $k$ times. So we can still obtain good independent classes, even when we have restricted the possible labelings k1k-1 times already. When pumping up a path for the first time, it results in a path of length $\Omega(n^{1/k})$, but when doing it for the second time, each node of that path might already have a pumped path of length $\Omega(n^{1/k})$ attached to it, so the resulting instance needs $\Omega(n^{2/k})$ many nodes. If we can still obtain an independent class in the graph, we show that the restrictions of the first indepndent classes are not too severe. Since we can still run our algorithm in an instance with roughly $n^{2/k}$ nodes, we obtain an independent class in the same way as above. If we tried to do this construction for more than $k$ steps, we would need more than $n$ nodes and so the simulation argument does not work anymore. Hence, an algorithm with complexity $o(n^{1/k})$ can be used to repeat the above process $k$ times. By then abusing the fact that this recursive decoupling can be done $k$ times, we obtain an algorithm with complexity $O(n^{1/(k+1)})$.
@@ 3 Locally Finite Labelings
Examples of MIS and 3-colouring as LFLs were already given in the introduction, but we additionally provide the problem of 3-cycle detection, as a more involved example in Appendix A.
@@
Additionally we provide one nontrivial extension to LFLs that we will use to make our proofs much more elegant. We call this extension partially ordered LFLs ($\mathbf{poLFL}$). In partially ordered LFLs we allow wildcard labels in the configurations. For example, we could then describe the problem of colouring using configurations where the center node has colour e.g. Red and it has a single optional neighbor with label Not Red. Now any neighbor that does not output colour Red can be mapped onto this optional neighbor. We illustrate how to encode 3 colouring in just 3 intuitive configurations this way in Figure 6. This formalism is not just nice in the sense, that it gives us much more freedom to express problems in a concise and intuitive way, but it also makes some of our more involved proofs more manageable.
Figure 6: Using the ”not __” labels ($\{\bar{R},\bar{G},\bar{B}\}$) we can encode 3-colouring in the very intuitive way, where a node that outputs $R$ must have only neighbors that are ”not RR\,”.
@@
3.1 LFLs in Trees - the Node-Edge Checkable Formalism
@@ A Locally Finite Labeling in the node–edge checkable formalism is a tuple
  • $\Sigma_{\text{out}}$ is a set of output labels, and we define $\Sigma_{\text{out}}^{*}=\{a,a^{*}\mid a\in\Sigma_{\text{out}}\}$. Thus for every output label $a$ we include a designated “optional” version $a^{*}$.
@@
Intuitively, each node $v$ inspects the multiset of labeled halfedges around it and checks whether some node constraint matches. If a pair $(a,b)$ appears exactly as often as required in a configuration $C$, the constraint is satisfied; if it appears more often, this is permitted only when $C$ contains $(a,b^{*})$, which serves as a the analog to an optional edge, allowing arbitrarily many occurrences.
MIS as a node-edge checkable problem
Using no input labels and the outputlabels $\left\{M,O,P\right\}$. The problem is given in the following way:
@@
Again this presentation of MIS is much more intuitive, then explicitly writing down all of these combinations for each possible $\Delta$.
@@ 3.2 The TwigLFL
Based on this observation, we add exactly $m_{\ell}(X)$ required edges leading to nodes with labeling $\ell$ to $C_{X}$. Additionally, we add an optional edge leading to a node labeled $\ell$ if and only if every configuration $C\in X$ with $m_{\ell}(C)=m_{\ell}(X)$ contains such an optional edge. By 9, this condition implies that in fact all configurations in $X$ contain such an optional edge.
@@
If $N$ contains more than $m_{\ell}(X)$ such nodes, then by 9 every configuration in $X$ must contain an optional edge labeled $\ell$, and hence $C_{X}$ contains exactly one such optional edge. We map all remaining $\ell$-labeled nodes of $N$ to this optional edge.